# A Backpropagation-Free Feedback–Hebbian Network for Continual Learning Dynamics

### v1

@@ Abstract
Feedback-rich neural architectures can regenerate earlier representations and inject temporal context, making them a natural setting for strictly local synaptic plasticity. We ask whether a minimal, backpropagation-free feedback–Hebbian system can already express interpretable continual-learning–relevant behaviors under controlled training schedules. We introduce a compact prediction–reconstruction architecture with two feedforward layers for supervised association learning and two dedicated feedback layers trained to reconstruct earlier activity and re-inject it as additive temporal context. All synapses are updated by a unified local rule combining centered Hebbian covariance, Oja-style stabilization, and a local supervised drive where targets are available, requiring no weight transport or global error backpropagation. On a small two-pair association task, we characterize learning through layer-wise activity snapshots, connectivity trajectories (row/column means of learned weights), and a normalized retention index across phases. Under sequential training, forward output connectivity exhibits a long-term depression (LTD)-like suppression of the earlier association while feedback connectivity preserves an -related trace during acquisition of . Under deterministic interleaving , both associations are concurrently maintained rather than sequentially suppressed. Architectural controls and rule-term ablations isolate the role of dedicated feedback in regeneration and co-maintenance, and the role of the local supervised term in output selectivity and unlearning. Together, the results show that a compact feedback pathway trained with local plasticity can support regeneration and continual-learning–relevant dynamics in a minimal, mechanistically transparent setting.
@@ I Introduction
Feedback-rich neural architectures are attractive because they can regenerate earlier representations, inject top–down context, and propagate temporal information that stabilizes and accelerates adaptation. Local update rules fit this regime: plasticity occurs where information flows, avoiding explicit global credit signals, with Hebbian mechanisms providing a simple and well-studied substrate for feedback-driven modulation and input regeneration. In cortex-like recurrent hierarchies, the interplay of feedforward drive, feedback, and priors supports generative prediction, while population analyses suggest that task-relevant information can occupy a stable mnemonic subspace despite rich dynamics; multiple circuit motifs (attractors, short-term synaptic processes, trained recurrent dynamics) can realize these functions [4, 7, 1].
@@
A substantial body of work explores alternatives to conventional weight transport and global error propagation. Feedback Alignment (FA) showed that deep networks can learn useful mappings even when backward pathways use fixed random weights [5]. Equilibrium Propagation (EqProp) casts learning in energy-based systems that use the same circuit in two phases, where a small output “nudge” yields a learning signal related to contrastive Hebbian learning and STDP [9]. Predictive coding introduces error-encoding units and can approximate supervised learning with local plasticity under certain conditions [10]. These directions primarily target credit assignment and biological plausibility; they are not typically framed around eliciting simple continual-learning (CL) primitives in a minimal feedback–Hebbian system. In contrast, mainstream CL methods are largely backpropagation-based and are commonly grouped into regularization, replay, and dynamic-architecture families [6]. Motivated by this gap, we ask whether a compact feedback–Hebbian model trained with strictly local plasticity can express interpretable CL-relevant behaviors under controlled protocols.
@@
We present a backpropagation-free feedback–Hebbian network with dedicated feedforward and feedback pathways trained by a strictly local plasticity rule that combines centered Hebbian covariance, Oja-style stabilization, and a local supervised drive where targets are available (Sec. II-C). Our objective is mechanistic: we study a minimal two-layer feedforward pathway paired with a two-layer feedback pathway on a small, controllable two-pair association task. Rather than emphasizing benchmark accuracy, we characterize learning by tracking layer-wise activity and weight evolution across training phases using simple connectivity summaries.
@@
Within this setting, we operationalize three continual-learning–relevant behaviors. (i) Retention under sequential training: after learning association A and then training only association B, A-related structure persists in the feedback pathway (as reflected by feedback-layer connectivity retention) even as the forward pathway adapts to express B at the output. (ii) LTD-like unlearning at the forward output stage: sequential AB training suppresses A-specific forward output connectivity while strengthening B-specific connectivity, yielding a targeted weakening reminiscent of synaptic depression. (iii) Co-maintenance under temporal alternation: under interleaved presentation, both associations are concurrently maintained, with task-relevant connectivity remaining elevated for both target sets rather than exhibiting sustained sequential suppression.
@@
Our contributions are threefold. (1) We introduce a minimal prediction–reconstruction architecture in which a dedicated feedback pathway is trained to reconstruct earlier-layer activity and is injected as additive temporal context. (2) We provide a unified local learning rule that is applied consistently to forward prediction and feedback reconstruction without weight transport or global error propagation. (3) Through layer-wise connectivity trajectories and phase-wise retention analyses, we show that this compact system expresses the defined continual-learning–relevant behaviors under sequential and interleaved training protocols.
@@ II-A Task and Data Representation
This dataset is inspired by cortical stimulation paradigms but used here to create a small, controllable association task. Inputs and targets are arrays corresponding to ten stimulation sites. Entries take values in ; in our experiments we use for clarity (0 = dormant, 1 = active), but the model does not require exact binary values.
We study two input positions (3, 7) and four target positions (5, 6, 8, 9) with fixed pairings: position 3 is linked to targets 8 and 9, and position 7 to targets 5 and 6. Concretely, when position 3 is set to 1 in the input, positions 8 and 9 are set to 1 in the target; when position 7 is set to 1 in the input, positions 5 and 6 are set to 1 in the target.
This yields a small multi-label mapping (multi-site input multi-site output) that is easy to probe and visualize. We use it to generate the phase schedules described in Sec. II-D, where we define the continual-learning diagnostics and their metrics, and to report the resulting weight/activation trajectories in Sec. III.
@@ II-B Model Structure
We instantiate the ten-site association task with a compact network that uses two forward layers and two dedicated feedback layers, each of size . The forward pathway computes the supervised inputoutput mapping in two stages; the feedback pathway provides a separate channel that reconstructs earlier activity and injects it as context at the next time step.
As shown in Fig. 1, green arrows denote forward computation and blue arrows denote feedback. Anatomical directions are fixed: short–term adaptation changes synaptic efficacy (weights), not wiring. Importantly, forward connections are not reused “in reverse”; backward signals travel on their own feedback connections. This keeps the roles distinct: forward layers carry prediction toward the outputs, while feedback layers reconstruct earlier activity to modulate subsequent processing via a lightweight, additive pathway.
The temporal use of feedback proceeds as follows. Let index discrete steps. At , no prior forward activity is available and feedback is inactive. For , the previous step’s activity in the later forward layer drives its paired feedback layer, which forms a reconstruction of the earlier layer’s state; this reconstruction is added to the earlier layer’s current drive. The merge is additive to preserve dimensionality and locality—no extra parameters or concatenation are introduced—yielding a lightweight local recurrence that carries recent context into the current computation.
We now justify the key structural choices, including depth, width, and the use of dedicated feedback. Exactly two forward layers constitute the minimal depth that separates the two roles we study here: the second forward layer supports supervised prediction to the outputs, and the associated feedback reconstructs the first forward layer’s activity to stabilize and guide subsequent inputs. The width of ten units per layer aligns with the ten stimulation sites, keeping parameter counts small and making weight and activity trajectories easy to inspect. Throughout, only the synaptic efficacies of these fixed connections are updated; all equations for the update rules and activations are given in the next subsection.
@@ II-C Algorithm
We consider a local plasticity rule that trains the learned connections in Fig. 1. Let denote the input to a trained layer and its output. Each trained weight matrix maps to (rows index output units; columns index input units). The same rule is applied to the two forward matrices and to the two dedicated feedback matrices. Supervision targets and which layers receive them are specified in Sec. II-D; for layers without supervision we set so the supervised term below is neutral.
@@
where is the learning rate, is the Oja-style normalization strength, and are the input, output, and (when present) target activities at epoch . Activities are centered by exponential running means
with an effective window of approximately . The first (covariance) term updates weights based on deviations from average activity, preventing units with persistently high activity from dominating irrespective of input statistics [3, 2]. The second (Oja) term implements activity-dependent normalization that counteracts runaway Hebbian growth [8]. The final supervised drive is local and biases updates toward the provided targets via ; when no supervision is provided for a layer, setting removes this contribution while leaving the other terms unchanged.
@@
which matches the dataset range, provides smooth bounded responses, and improves stability for local Hebbian updates in the presence of feedback. Unbounded activations such as ReLU can amplify Hebbian growth under feedback unless additional controls are introduced.
Unless noted otherwise, we use , , and . Weights are initialized with a standard variance-scaled uniform distribution using fan-in scaling. Running means and are initialized to zero and carried across protocol phases to reflect ongoing baseline drift.
These design choices are motivated as follows. Centering directs plasticity to deviations from the current baseline, preventing persistently active units from dominating irrespective of input statistics and mitigating interference across phases. The Oja term constrains row norms and induces competition, enabling activity-dependent depression when output activity is high without matching input drive; this both prevents runaway growth and supports equalization during unlearning phases. The supervised drive accelerates acquisition while maintaining locality and leaving normalization intact. The bounded activation caps signal magnitudes, stabilizing dynamics when feedback injects recent activity. Empirical effects on retention, conditioning, and equalization are reported in Sec. III. Updates are local and per synapse with cost per step; no gradients or replay buffers are stored.
@@ II-D Training Procedure
Training uses a local, backpropagation-free rule (Eq. (1)) to optimize forward prediction and feedback reconstruction. No weight transport is used; all plasticity is local.
Within each epoch, activations follow Sec. II-C. Supervision is applied only where targets exist: forward layer 2 is driven toward the task output vector, and each feedback layer is driven to reconstruct the input to its paired forward layer. Forward layer 1 receives no explicit target; its weights adapt locally from the incoming activity and Oja normalization. Equation (1) is applied once to each learned matrix with additive updates using activities centered by exponential running means. The running means and persist across phases.
@@
Regime I: Sequential A then B (LTD-like effect and retention). Training proceeds in two phases. Phase A presents only pair A samples (input site 3 mapping to output sites 8 and 9), 50 samples per epoch for 10 epochs. Phase B then presents only pair B samples (input site 7 mapping to output sites 5 and 6), 50 samples per epoch for 10 epochs. Pair A is not presented during Phase B. This protocol tests depression of the previously learned association in the forward pathway during B-training, an LTD-like effect under the local rule, and retention of A-related connectivity in the feedback pathway.
@@ II-E Connectivity and Retention Metrics
Each trained synaptic matrix has rows indexing output units and columns indexing input units. For a given matrix, input connectivity is defined as the column-wise mean, yielding a profile in which each entry reflects how strongly the corresponding input site projects into that layer on average. Conversely, output connectivity is defined as the row-wise mean, yielding a profile in which each entry reflects how strongly the corresponding output unit integrates its inputs on average. These input and output connectivity profiles are computed separately for each trained forward and feedback matrix for all training protocols. The resulting trajectories are visualized in Sec. III to illustrate how synaptic strengths redistribute across sites and layers over time.
In addition to these trajectories, we use a normalized retention index to summarize how strongly a given connectivity measure changes during a later training phase, relative to the magnitude of the change induced by an earlier phase. Let denote a scalar connectivity summary at epoch for a chosen layer. The specific choice of for a given analysis is described in Sec. III when reporting retention values; the definition of the normalized index itself is independent of that choice.
@@
When retention values are reported in Sec. III, the corresponding baseline, pre-phase, and post-phase epochs, as well as the underlying choice of connectivity summary , are specified with the description of each protocol.
@@ III Results
We evaluate whether the feedback–Hebbian network exhibits three continual–learning–relevant primitives on the two-pair association task introduced in Sec. II. These primitives are inspired by circuit-level phenomena in neuroscience, but are instantiated here as operational behaviors in a small, controlled supervised setting.
@@
Our analyses proceed in two stages. We first verify baseline activity-level behavior (Sec. III-A) by checking that output activation concentrates on target sites and that feedback regenerates the corresponding input pattern from the output. We then interpret sequential and interleaved learning using the connectivity profiles and normalized retention index defined in Sec. II, which summarize how task-relevant synaptic strengths evolve from baseline through earlier and later training phases. Subsequent subsections apply this framework to single-association learning and regeneration, sequential AB training, interleaved A/B conditioning, architectural controls, and rule-term ablations.
@@ III-A Single-Association Learning and Regeneration
We first verify the two baseline capabilities of the proposed 2F+2B network: (i) learning a single input–output association in the forward direction and (ii) regenerating the corresponding input pattern via the dedicated feedback pathway. The network is trained on pair A only (Phase A of Regime I; Sec. II-D).
Forward prediction.
In the Pre-training snapshot (random initialization), the output response to the pair A input is unstructured (Fig. 2). After completing Phase A training (Post Phase A snapshot), the same input elicits a selective output in which the dominant components coincide with the pair A target sites (outputs 8 and 9), while non-target sites are driven toward zero (Fig. 2).
Feedback-driven regeneration.
We next probe whether the feedback pathway can regenerate the input-side pattern associated with pair A. When the feedback pathway is driven by the pair A output pattern, the reconstructed activity is diffuse in the Pre-training snapshot (Fig. 3). After Phase A training (Post Phase A snapshot), the reconstruction is dominated by the originally stimulated input site (site 3), with non-target sites driven toward zero (Fig. 3).

### v2

@@ Abstract
Feedback-rich neural architectures can regenerate earlier representations and inject temporal context, making them a natural setting for strictly local synaptic plasticity. Existing literature raises doubt about whether a minimal, backpropagation-free feedback–Hebbian system can already express interpretable continual-learning–relevant behaviors under controlled training schedules. In this work, we introduce a compact prediction–reconstruction architecture with a dedicated feedback pathway providing lightweight, locally trainable temporal context for continual adaptation. All synapses are updated by a unified local rule combining centered Hebbian covariance, Oja-style stabilization, and a local supervised drive where targets are available. With a simple two-pair association task, learning is characterized through layer-wise activity snapshots, connectivity trajectories (row/column means of learned weights), and a normalized retention index across phases. Under sequential training, forward output connectivity exhibits a long-term depression (LTD)-like suppression of the earlier association while feedback connectivity preserves an -related trace during acquisition of . Under alternating sequence, both associations are concurrently maintained rather than sequentially suppressed. Architectural controls and rule-term ablations isolate the role of dedicated feedback in regeneration and co-maintenance alongside the role of the local supervised term in output selectivity and unlearning. Together, the results show that a compact feedback pathway trained with local plasticity can support regeneration and continual-learning–relevant dynamics in a minimal, mechanistically transparent setting.
@@ I Introduction
Feedback-rich neural architectures are attractive because they can regenerate earlier representations, inject top–down context, and propagate temporal information that stabilizes and accelerates adaptation [7]. Local update rules are suitable to this regime: plasticity occurs where information flows, avoiding explicit global credit signals, with Hebbian mechanisms providing a simple and well-studied substrate for feedback-driven modulation and input regeneration. In cortex-like recurrent hierarchies, the interplay of feedforward drive, feedback, and priors supports generative prediction, while population analyses suggest that task-relevant information can occupy a stable mnemonic subspace despite rich dynamics, where multiple circuit motifs (attractors, short-term synaptic processes, trained recurrent dynamics) can realize these functions [1, 10].
@@
A substantial body of work explores alternatives to conventional weight transport and global error propagation. Feedback Alignment (FA) showed that deep networks can learn useful mappings even when backward pathways use fixed random weights [8]. Equilibrium Propagation (EqProp) casts learning in energy-based systems that use the same circuit in two phases, where a small output “nudge” yields a learning signal related to contrastive Hebbian learning and Spike-timing-dependent plasticity (STDP) [13]. Predictive coding introduces error-encoding units and can approximate supervised learning with local plasticity under certain conditions [14]. These directions primarily target credit assignment and biological plausibility; they are not typically framed around eliciting simple continual-learning (CL) primitives in a minimal feedback–Hebbian system. In contrast, mainstream CL methods are largely backpropagation-based and are commonly grouped into regularization, replay, and dynamic-architecture families [9]. Motivated by this gap, we raise a doubt about whether a compact feedback–Hebbian model trained with strictly local plasticity can express interpretable CL-relevant behaviors under controlled protocols.
@@
This study proposes a backpropagation-free feedback–Hebbian network with dedicated feedforward and feedback pathways trained by a strictly local plasticity rule that combines centered Hebbian covariance, Oja-style stabilization, and a local supervised drive where targets are available. Our objective is mechanistic: we study a minimal two-layer feedforward pathway paired with a two-layer feedback pathway on a small, controllable two-pair association task. Rather than emphasizing benchmark accuracy, we characterize learning by tracking layer-wise activity and weight evolution across training phases using simple connectivity summaries.
@@
Within this setting, we operationalize three continual-learning–relevant behaviors. (i) Retention under sequential training: after learning association A and then training with association B only, A-related structure persists in the feedback pathway (as reflected by feedback-layer connectivity retention) even as the forward pathway adapts to express B at the output. (ii) LTD-like unlearning at the forward output stage: sequential AB training suppresses A-specific forward output connectivity while strengthening B-specific connectivity, yielding a targeted weakening reminiscent of synaptic depression. (iii) Co-maintenance under temporal alternation: under alternating sequence (e.g., ), both associations are concurrently maintained, with task-relevant connectivity remaining elevated for both target sets rather than exhibiting sustained sequential suppression.
@@
Our contributions include: (1) We introduce a minimal prediction–reconstruction architecture in which a dedicated feedback pathway is trained to reconstruct earlier-layer activity and is injected as additive temporal context. (2) We provide a unified local learning rule that is applied consistently to forward prediction and feedback reconstruction without weight transport or global error propagation. (3) Through layer-wise connectivity trajectories and phase-wise retention analyses, we show that this compact system expresses the defined continual-learning–relevant behaviors under sequential and interleaved training protocols.
@@ II-A Task and Data Representation
Our dataset is inspired by cortical stimulation paradigms but used here to create a small, controllable association task. Inputs and targets are arrays corresponding to ten stimulation sites. Entries take values in ; in our experiments we use for clarity (0 = dormant, 1 = active), but the model does not require exact binary values.
We study two input positions (3, 7) and four target positions (5, 6, 8, 9) with fixed pairings: Pair A has position 3 that links to targets 8 and 9, and Pair B has position 7 links to targets 5 and 6. Concretely, when position 3 is set to 1 in the input, positions 8 and 9 are set to 1 in the target; when position 7 is set to 1 in the input, positions 5 and 6 are set to 1 in the target.
This yields a small multi-label mapping (multi-site input multi-site output) that is easy to probe and visualize. We use it to generate the phase schedules, where we define the continual-learning diagnostics and their metrics, and to report the resulting weight/activation trajectories.
@@ II-B Model Structure
We instantiate the ten-site association task with a compact network that uses two forward layers and two dedicated feedback layers, each of size . The forward pathway computes the supervised inputoutput mapping in two stages, and the feedback pathway provides a separate channel that reconstructs earlier activity and injects it as context at the next time step.
As shown in Fig. 1, green arrows denote forward computation and blue arrows denote feedback. Anatomical directions are fixed: short–term adaptation changes synaptic efficacy (weights), not wiring. More importantly, forward connections are not reused “in reverse”, as backward signals travel on their own feedback connections. This keeps the roles distinct: forward layers carry prediction toward the outputs, while feedback layers reconstruct earlier activity to modulate subsequent processing via a lightweight, additive pathway.
The temporal use of feedback proceeds as follows. Let index discrete steps. At , no prior forward activity is available and feedback is inactive. For , the previous step’s activity in the forward layer drives its paired feedback layer, forming a reconstruction of the earlier layer’s state, which is then added to the forward layer’s current drive. The merge is additive to preserve dimensionality and locality—no extra parameters or concatenation are introduced—yielding a lightweight local recurrence that carries recent context into the current computation.
We now justify the key structural choices, including depth, width, and the use of dedicated feedback. Exactly two forward layers constitute the minimal depth that separates the two roles we study here: the second forward layer supports supervised prediction to the outputs, and the associated feedback reconstructs the first forward layer’s activity to stabilize and guide subsequent inputs. The width of ten units per layer aligns with the ten stimulation sites, keeping parameter counts small and making weight and activity trajectories easy to inspect. Therefore, only the synaptic efficacies of these fixed connections are updated.
@@ II-C Algorithm
We consider a local plasticity rule that trains the learned connections in Fig. 1. Let denote the input to a trained layer and its output. Each trained weight matrix maps to (rows index output units; columns index input units). The same rule is applied to the two forward matrices and to the two dedicated feedback matrices. For layers without supervision we set such that the supervised term is neutral.
@@
where is the learning rate, is the Oja-style normalization strength, and are the input, output, and (when present) target activities at epoch respectively. Activities are centered by exponential running means
where is a constant smoothing parameter that sets the effective averaging window to approximately . The first (covariance) term updates weights based on deviations from average activity, preventing units with persistently high activity from dominating irrespective of input statistics [6, 2]. The second (Oja) term implements activity-dependent normalization that counteracts runaway Hebbian growth [11]. The final supervised drive is local and biases updates toward the provided targets via ; when no supervision is provided for a layer, setting removes this contribution while leaving the other terms unchanged.
@@
where denotes the layer output pre-activation. The function matches the dataset range, provides smooth bounded responses, and improves stability for local Hebbian updates in the presence of feedback. Unbounded activations such as ReLU can amplify Hebbian growth under feedback unless additional controls are introduced.
Unless noted otherwise, we use , , and . Weights are initialized with a standard variance-scaled uniform distribution using fan-in scaling [5]. Running means and are initialized to zero and carried across protocol phases to reflect ongoing baseline drift.
Centering directs plasticity to deviations from the current baseline, preventing persistently active units from dominating irrespective of input statistics and mitigating interference across phases. The Oja term constrains row norms and induces competition, enabling activity-dependent depression when output activity is high without matching input drive; this both prevents runaway growth and supports equalization during unlearning phases. The supervised drive accelerates acquisition while maintaining locality and leaving normalization intact. The bounded activation caps signal magnitudes, stabilizing dynamics when feedback injects recent activity. Updates are local and per synapse with cost per step; no gradients or replay buffers are stored.
@@ II-D Training Procedure
Training uses the aforementioned local, backpropagation-free rule (Eq. (1)) to optimize forward prediction and feedback reconstruction. Weight transport is unnecessary, while all plasticity is local.
Supervision is applied only where targets exist: forward layer 2 is driven toward the task output vector, and each feedback layer is driven to reconstruct the input to its paired forward layer. Forward layer 1 receives no explicit target; its weights adapt locally from the incoming activity and Oja normalization. Eq. (1) is applied once to each learned matrix with additive updates using activities centered by exponential running means. The running means and persist across phases.
@@
Regime I: Sequential A then B (LTD-like effect and retention). Training proceeds in two phases. Phase 1 presents only pair A samples (input site 3 mapping to output sites 8 and 9), 50 samples per epoch for 10 epochs. Phase 2 then presents only pair B samples (input site 7 mapping to output sites 5 and 6), 50 samples per epoch for 10 epochs. Pair A is not presented during Phase 2. This protocol tests depression of the previously learned association in the forward pathway during Phase 2, an LTD-like effect under the local rule, and retention of A-related connectivity in the feedback pathway.
@@ II-E Connectivity and Retention Metrics
Each trained synaptic matrix has rows indexing output units and columns indexing input units. For a given matrix, input connectivity is defined as the column-wise mean, yielding a profile in which each entry reflects how strongly the corresponding input site projects into that layer on average. Conversely, output connectivity is defined as the row-wise mean, yielding a profile in which each entry reflects how strongly the corresponding output unit integrates its inputs on average. These input and output connectivity profiles are computed separately for each trained forward and feedback matrix for all training protocols. The resulting trajectories are then visualized to illustrate how synaptic strengths redistribute across sites and layers over time.
In addition to these trajectories, we use a normalized retention index to summarize how strongly a given connectivity measure changes during a later training phase, relative to the magnitude of the change induced by an earlier phase. Let denote a scalar connectivity summary at epoch for a chosen layer.
@@
When retention values are reported, the corresponding baseline, pre-phase, and post-phase epochs, as well as the underlying choice of connectivity summary , are specified alongside each protocol’s results.
@@ III Results
in this experiment, we evaluate whether the feedback–Hebbian network exhibits three continual–learning–relevant primitives on the two-pair association task. These primitives are inspired by circuit-level phenomena in neuroscience [4, 12, 3], but are instantiated here as operational behaviors in a small, controlled supervised setting.
@@
Our analyses proceed in two stages. We first verify baseline activity-level behavior by checking that output activation concentrates on target sites and that feedback regenerates the corresponding input pattern from the output. We then interpret sequential and interleaved learning using the connectivity profiles and normalized retention index, which summarize how task-relevant synaptic strengths evolve from baseline through earlier and later training phases. Here, we apply this framework to single-association learning and regeneration, sequential AB training, interleaved A/B conditioning, architectural controls, and rule-term ablations.
@@ III-A Single-Association Learning and Regeneration
We first verify the two baseline capabilities of the proposed 2FF+2FB network: (i) learning a single input–output association in the forward direction and (ii) regenerating the corresponding input pattern via the dedicated feedback pathway. The network is trained on pair A only (Phase 1 of the sequential A-then-B training regime).
Forward prediction
In the Pre-training snapshot (random initialization), the output response to the pair A input is unstructured, as shown in Fig. 2. After completing Phase 1 training (Post Phase 1 snapshot), the same input elicits a selective output in which the dominant components coincide with the pair A target sites (outputs 8 and 9), while non-target sites are driven toward zero.
Feedback-driven regeneration
We then probe whether the feedback pathway can regenerate the input-side pattern associated with pair A. When the feedback pathway is driven by the pair A output pattern, the reconstructed activity is diffuse in the Pre-training snapshot as shown in Fig. 3. After Phase 1 training (Post Phase 1 snapshot), the reconstruction is dominated by the originally stimulated input site 3, with non-target sites driven toward zero.